### v2

# A survey and benchmark of high-dimensional Bayesian optimization of discrete sequences

Abstract
Optimizing discrete black box functions is key in several domains, e.g. protein engineering and drug design. Due to the lack of gradient information and the need for sample efficiency, Bayesian optimization is an ideal candidate for these tasks. Several methods for high-dimensional continuous and categorical Bayesian optimization have been proposed recently. However, our survey of the field reveals highly heterogeneous experimental set-ups across methods and technical barriers for the replicability and application of published algorithms to real-world tasks. To address these issues, we develop a unified framework to test a vast array of high-dimensional Bayesian optimization methods and a collection of standardized black box functions representing real-world application domains in chemistry and biology. These two components of the benchmark are each supported by flexible, scalable, and easily extendable software libraries (poli and poli-baselines), allowing practitioners to readily incorporate new optimization objectives or discrete optimizers. Project website: https://machinelearninglifescience.github.io/hdbo_benchmark.
1 Introduction
Optimizing an unknown and expensive-to-evaluate function is a frequent problem across disciplines (Shahriari et al.,, 2016), examples are finding the right parameters for machine learning models or simulators, drug discovery (Gómez-Bombarelli et al.,, 2018; Griffiths and Hernández-Lobato,, 2020; Pyzer-Knapp,, 2018), protein design (Stanton et al.,, 2022; Gruver et al.,, 2023), hyperparameter tuning in Machine Learning (Snoek et al.,, 2012; Turner et al.,, 2021) and train scheduling. In some scenarios, evaluating the black box involves an expensive process (e.g. training a large model, or running a physical simulation); Bayesian Optimization (BO, Močkus, (1975)) is a powerful method for sample efficient black box optimization. High dimensional (discrete) problems have long been identified as a key challenge for Bayesian optimization algorithms (Wang et al.,, 2013; Snoek et al.,, 2012) given that they tend to scale poorly with both dataset size and dimensionality of the input.
High-dimensional BO has been the focus of an entire research field (see Fig. 1), in which methods are extended to address the curse of dimensionality and its consequences (Binois and Wycoff,, 2022; Santoni et al.,, 2023). Within this setting, discrete sequence optimization has received particular focus, due to its applicability in the optimization of molecules and proteins. However, prior work often focuses on sequence lengths and number of categories below the hundreds (see Fig. 2), making it difficult for practitioners to judge expected performance on real-world problems in these domains. We contribute (i) a survey of the field while focusing on the real-world applications of high-dimensional discrete sequences, (ii) a benchmark of several optimizers on established black boxes, and (iii) an open source, unified interface: poli and poli-baselines.111https://github.com/MachineLearningLifeScience/{poli, poli-baselines}
2 Preliminaries
Bayesian Optimization and Gaussian processes.
Bayesian optimization requires a surrogate model and an acquisition function (Garnett,, 2023). Given both, the objective function is sequentially optimized by fitting a model to the given observations and numerically optimizing the acquisition function with respect to the model to select the next configuration for evaluation. Frequently, the model is a Gaussian process (GP, Rasmussen and Williams, (2006)), and popular choices for the acquisition function are Expected Improvement (Jones et al.,, 1998; Garnett,, 2023) and the Upper Confidence Bound (Srinivas et al.,, 2012). A GP allows to express a prior belief over functions. Formally, it is a collection of random variables, such that every finite subset follows a multivariate normal distribution, described by a mean function , and a positive definite covariance function (kernel) (Rasmussen and Williams,, 2006, p. 13). Assuming that observations of the function are distorted by Gaussian noise, the posterior over the function conditioned on these observations is again Gaussian. The prediction equations have a closed form and can be evaluated in time where is the number of observations.
Is high-dimensional Bayesian Optimization difficult?
There are three reasons why BO is thought to scale poorly with dimension. The first reason is that GPs fail to properly fit to the underlying objective function in high dimensions. Secondly, even if the GPs were to fit well there is still the problem of optimizing the high-dimensional acquisition function. Finally, Gaussian Processes are believed to scale poorly with the size of the dataset, limiting us to low-budget scenarios (Binois and Wycoff,, 2022). Folk knowledge suggests that GPs fail to fit functions above the meager limit of dimensions (Santoni et al.,, 2023) and datapoints.
Hvarfner et al., (2024) recently disputed these well-entrenched narratives by showing that poor GP fitting could be caused by a poor choice of regularizer; mitigating the curse of dimensionality could be as easy as including a dimensionality-dependent prior over lengthscales. Furthermore, Xu and Zhe, (2024) argues that even the simplest BO outperforms highly elaborate methods.
Optimization of discrete sequences & applications.
Most HDBO methods are tested on toy examples, hyperparameter tuning, or reinforcement learning tasks (Binois and Wycoff,, 2022; Penubothula et al.,, 2021). We focus on discrete sequence optimization, which has several applications beyond the usual examples (e.g. MaxSAT, or PestControl) (Papenmeier et al.,, 2024), and is key in applications to biology and bioinformatics (Gómez-Bombarelli et al.,, 2018; Stanton et al.,, 2022; Gruver et al.,, 2023). Drug design and protein engineering can be thought of as sequence optimization problems, if we consider the SMILES/SELFIES representation of small molecules (Weininger,, 1988; Krenn et al.,, 2020), or the amino acid sequence representation of proteins (Needleman and Wunsch,, 1970).
Related work.
Binois and Wycoff, (2022) initially surveyed the field of high-dimensional GPs, focusing on applications to BO, and proposed a taxonomy of structural assumptions for GPs that includes variable selection, additive models, linear, and non-linear embeddings. This work has since been updated by Wang et al., (2023) and Santoni et al., (2023). The latter presents an empirical study on the continuous, toy-problem setting up to 60 dimensions and refines the taxonomy (Binois and Wycoff,, 2022) into five categories, separating trust regions from the rest. Griffiths et al., (2023) compare different kernel functions used with binary vector representations of molecules and for the same application Kristiadi et al., (2024) study the use of different large language models. Our work is most similar to Dreczkowski et al., (2023)’s comprehensive overview of discrete BO (MCBO), and Gao et al., (2022)’s benchmark of small molecule optimization. In both, HDBO is not in focus.
3 A taxonomy of high-dimensional Bayesian Optimization
We describe the field of high dimensional BO and the large number of related publications through a refined taxonomy building on previous work, discussing variable selection, additive models, trust regions, linear embeddings, non-linear embeddings, gradient information, structured spaces, and others in turn. While encompassing taxonomies over fields may initially appear ill-advised (Wilkins,, 1668, pp.22), we highlight commonalities in strategies that give structure to the HDBO problem-space.
We expand previous surveys (Binois and Wycoff,, 2022; Santoni et al.,, 2023) and identify a finer taxonomy of seven method groups and new families of structured spaces (i.e. methods that work directly on mixed representations, or Riemannian manifolds, previously categorized as non-linear embeddings), and methods that rely on predicted gradient information. This new separation emphasizes the heterogeneous nature of discrete solvers: some optimizers work directly on discrete space (structured spaces), while others optimize using latent representations (non-linear embeddings); gradient-based methods are separated to show alternatives when first-order information is available or modelable. Fig. 1 presents a timeline of HDBO methods, split into these families, and all methods are detailed in supplementary Table LABEL:tab:appendix:full_taxonomy; methods are grouped according to their most dominant feature.
3.1 Variable selection
To solve a high-dimensional problem, one approach is to focus on a subset of variables of high interest.222Under the assumption that there exists an axis-aligned lower-dimensional active subspace. One selects the variables either by using domain expertise, or by Automatic Relevance Detection (ARD) (Rasmussen and Williams,, 2006, pp.106-107) i.e. large lengthscales indicate independence under the covariance matrix for GPs. Examples of this approach include Hierarchical Diagonal Sampling (HDS) (Chen et al.,, 2012) and the Dimension Scheduling Algorithm (DSA) (Ulmasov et al.,, 2016). The former determines the active variables by a binary tree of subsets of , and fits GPs in lower-dimensional projections. DSA constructs a probability distribution by the principal directions of the training inputs and subsamples the dimensions accordingly. In contrast Li et al., (2018) randomly sample subsets of active dimensions.
Other methods rely on placing priors on their lengthscales, followed by a Bayesian treatment of the training. In Sequential Optimization of Locally Important Directions (SOLID), lengthscales are weighted by a Bernoulli distributed parameter, and coordinates are removed when their posterior probability goes below a user-specified threshold. (Winkel et al.,, 2021). Eriksson and Jankowiak, (2021) consider the Sparse Axis-Aligned Subspace (SAAS) model of a GP, restricting the function space through a (long-tailed) half-Cauchy prior on the inverse-lengthscales of the kernel.
3.2 Additive models
Additive models assume that the objective function can be decomposed into a sum of lower-dimensional functions. Symbolically, the coordinates of a given input are split into usually disjoint subgroups of smaller size, called a decomposition. Instead of fitting a GP to variables in , the algorithm fits GPs to the restrictions and adds their Upper Confidence Bound. The differences between the algorithms in this family are on how the subgroups are constructed, how the additive structure is approximated, the training of the Gaussian Process, or leveraging special features (Mutny and Krause,, 2018).
Han et al., (2021) select the decomposition which maximizes the marginal likelihood from a collection of randomly sampled decompositions, updating it every certain number of iterations. Alternatives include: leveraging a generalization based on restricted projections (Li et al.,, 2016), discovering the additive structure using model selection and Markov Chain Monte Carlo (Gardner et al.,, 2017), considering overlapping groups (Rolland et al.,, 2018), ensembles of Mondrian space-tiling trees (Wang et al.,, 2018), or use random tree-based decompositions (Ziomek and Bou Ammar,, 2023).
3.3 Trust regions
Some BO algorithms restrict the evaluation of the acquisition function to a small region of input space called a trust region, which is centered at the incumbent and is dynamically contracted or expanded according to performance (Regis,, 2016; Pedrielli and Ng,, 2016; Eriksson et al.,, 2019). Contemporary variants extend to the multivariate setting (e.g. MORBO (Daulton et al., 2022a, )), to quality-diversity (Maus et al.,, 2023) and to the optimization of mixed variables (CASMOPOLITAN by Wan et al., (2021)), including categorical. Since the trust region framework involves only the optimization of the acquisition function, several other methods leverage it alongside other structural assumptions like linear/non-linear embeddings (e.g. Tripp et al., (2020); Papenmeier et al., (2022)).
3.4 Linear embeddings
Instead of optimizing directly in input space , several methods rely on optimizing in a lower-dimensional space , which is linearly embedded into data space using a linear transformation (Wang et al.,, 2016). The matrix can be either selected at random (Wang et al.,, 2016; Qian et al.,, 2016), computed as a low-rank approximation of the input data matrix (Djolonga et al.,, 2013; Zhang et al.,, 2019; Raponi et al.,, 2020), constructed using gradient information and active subspaces (Palar and Shimoyama,, 2017; Wycoff et al.,, 2021), or through the minimization of variance estimates (Hu et al.,, 2024).
These methods are limited by how low-dimensional exploration translates into high dimensions. One choice of embedding matrix spans a fixed, highly-restricted subspace of . For this approach several issues regarding back-projections need to be addressed. Indeed, projecting from bounded domains to might render points outside the bounded domain in the input (Binois and Wycoff,, 2022). Finally, the transformation is not injective, meaning a point in input space can correspond to several latent points (Binois et al.,, 2015; Moriconi et al., 2020b, ).
Binois et al., (2015) propose a kernel that alleviates these issues by including a back-projection to the bounded domain that respects distances in the embedded space. Hashing matrices are an alternative way to reconstruct an input data point in a bounded domain from a latent point , whose entries are either 0, 1, and -1. Thus, the result of multiplying is a linear combination of the coordinates of where the coefficients are 1 and -1 (Nayebi et al.,, 2019). These ideas have been combined with trust regions both in the continuous (Papenmeier et al.,, 2022) and mixed-variable settings (Papenmeier et al.,, 2024). A natural extension considers a family of nested subspaces, progressively growing the embedding matrix until it matches the input dimensionality (Papenmeier et al.,, 2022). An alternative that does not deal with reconstruction mappings (thus circumventing the aforementioned issues) uses the information learned in the lower dimensional space to perform optimization directly in input space (Horiguchi et al.,, 2022).
3.5 Non-linear embeddings
Several methods have considered non-linear embeddings to incorporate learned latent representations. One set of examples are deep latent variable models like Generative Adversarial Networks (Goodfellow et al.,, 2014), or variants of Autoencoders (Kingma and Welling,, 2014; Stanton et al.,, 2022; Maus et al.,, 2022), algorithms that allow for modelling arbitrarily structured inputs. This is highly relevant for optimizing sequences, which are modeled as samples from a categorical distribution.
Gómez-Bombarelli et al., (2018) pioneered latent space optimization (LSBO) by learning a latent space of small molecules through their SMILES representation using a Variational Autoencoder (VAE, Kingma and Welling, (2014); Rezende et al., (2014)), and optimizing metrics such as the qualitative estimate of druglikeness (QED) therein. Several approaches have followed, including usage of a-priori given labelled data (Eissman et al.,, 2018) or decoder uncertainty (Notin et al.,, 2021), smart retraining schemes that focus on promising points (Tripp et al.,, 2020), metric-learning approaches that match promising points together (Grosnit et al.,, 2021), constraining the latent space (Griffiths and Hernández-Lobato,, 2020), latent spaces mapping to graphs (Kusner et al.,, 2017; Jin et al.,, 2018) and jointly learning the surrogate model and the latent representation (Maus et al.,, 2022; Lee et al.,, 2023; Chen et al.,, 2024; Kong et al.,, 2024). Stanton et al., (2022) take this further by learning multiple representations: one shared and required for both the decoder and surrogate, and one discriminative encoding as input for a GP used in the acquisition function. A prerequisite for these methods is a large dataset of unsupervised inputs, which may not be available in all applications. The methods that rely on training both the representation and the regression at the same time need supervised labels, which may be potentially unavailable. Optimization in embedding spaces greatly increases the complexity of problems that can be tackled, making it an appealing alternative for discrete sequence optimization in real-world tasks (see Fig. 2).
3.6 Gradient information
High-dimensional problems can become significantly easier when derivative information is available. Even when the objective’s derivatives are not available, the gradient information from the surrogate model can guide exploration. In our case, the referenced approaches cannot be applied directly, as they assume a differentiable kernel. For methods that rely on a continuous latent representation (see Secs. 3.4 and 3.5), gradient information of the surrogate model in latent space can be used.
Ahmed et al., (2016) mention how several Bayesian optimization methods could leverage gradient information and encourage the community to augment their optimization schemes with gradients, supported by strong empirical results even with randomly sampled directional derivatives. Eriksson et al., (2018) alleviate the computational constraints that come from using supervised gradient information using structured kernel interpolation and computational tricks like fast matrix-vector multiplication and pivoted Cholesky preconditioning. Other avenues for mitigating the computational complexity involve using structured automatic differentiation (Ament and Gomes,, 2022). Instead of using the gradient for taking stochastic steps, Penubothula et al., (2021) aim to find local critical points by querying where the predicted gradient is zero.
As mentioned above, fitting a Gaussian process to the objective allows for predicting gradients without having seen them a priori (Rasmussen and Williams,, 2006, Sec 9.4); Müller et al., (2021) propose Gradient Information with BO (GIBO), in which they guide local policy search in reinforcement learning tasks, exploiting this property. Nguyen et al., (2022) address that expected gradients may not lead to the best performing outputs and compute the most probable descent direction.
3.7 Structured spaces
Some applications work over structured spaces. For example, the angles of robot arms and protein backbones map to Riemannian manifolds (Jaquier et al.,, 2020; Penner,, 2022), and input spaces might also contain mixed variables (i.e. products of real and categorical spaces). To compute non-linear embeddings (see Sec. 3.5) followed by standard Bayesian optimization (or small variations thereof) can allow us to work over such spaces. Jaquier et al., (2020) use kernels defined on Riemannian manifolds (Feragen et al.,, 2015; Borovitskiy et al.,, 2020) and optimize the acquisition function using tools from Riemannian optimization (Boumal,, 2023). The authors expand their framework to high-dimensional manifolds by projecting to lower-dimensional submanifolds, which is roughly the equivalent to linear embeddings in the Riemannian settings (Jaquier and Rozo,, 2020).
In the categorical and mixed-variable setting, kernels over string spaces (Lodhi et al.,, 2000; Shervashidze et al.,, 2011), can be applied to BO (Moss et al.,, 2020). Other methods construct combinatorial graph and diffusion kernels-based GPs (Oh et al.,, 2019). Deshwal and Doppa, (2021) combine latent space kernels with combinatorial kernels in an autoencoder-based approach.
Recently, Daulton et al., 2022b have proposed a continuous relaxation of the discrete variables to ease the optimization of the acquisition function. Deshwal et al., (2023) propose another way to map discrete variables to continuous space, relying on Hamming distances to make a dictionary for embeddings. Papenmeier et al., (2024) extend previous work to both continuous and categorical variables: BAxUS learns an increasing sequence of subspaces using hash matrices which, when combined with the CoCaBo kernel (Ru et al.,, 2020), renders an algorithm for the mixed-variable setting. Finally, through a continuous relaxation of the objective that incorporates prior pretrained models, Michael et al., (2024) propose a surrogate on the probability vector space to optimize either the discrete input space or a continuous latent one.
Other.
Some methods evade our taxonomy but are worth mentioning: some focus on the optimization of the acquisition function and the impact of initializations (Zhao et al.,, 2024; Ngo et al.,, 2024). Other methods balance both active learning (i.e. building a better surrogate model) and optimization (Hvarfner et al.,, 2023). Most recently, two articles claimed that the standard setting for Bayesian optimization or slight variations of it perform as well as the state-of-the-art of all the aforementioned families (Hvarfner et al.,, 2024; Xu and Zhe,, 2024) – begging the question, can these methods optimize in high dimensional discrete problem spaces in a sample efficient manner?
4 Benchmarking the performance of HDBO methods
Practitioners that decide on what Bayesian optimization algorithm to use for their application will face several challenges. While surveying the field, we noticed two key discrepancies in the reported experimental set-ups: (i) the initialization varies from as low as none randomly/SOBOL sampled points to over , (ii) evaluation budgets also vary for the same types of tasks. Fig. 3 visualizes these different experimental set-ups as swarmplots. Moreover, our survey covered code availability. The state-of-the-art is being pulled by workhorses, which have democratized access to GP and BO implementations: GPyTorch (Gardner et al.,, 2018) and BoTorch (Balandat et al.,, 2020), and GPFlow (Matthews et al.,, 2017; van der Wilk et al.,, 2020) and Trieste (Picheny et al.,, 2023). These libraries are highly useful and impactful, yet one can obtain cross-dependency conflicts between them especially if third-party dependencies are introduced or if very specific versions are required for solver setups. As a particular example, solvers like ProbRep cannot co-exist with Ax-based solvers like SAASBO or Hvarfner’s VanillaBO. There is a need for isolating optimizers, specifying up-to-date environments in which they can run. These issues led to the development of poli.
4.1 poli and poli-baselines: a framework for benchmarking discrete optimizers
We want to solve truly high dimensional problems that are relevant for domains like biology and chemistry. To make the outcomes comparable, we require a unified way of defining the problem which includes consistent starting points, budgets, runtime environments, relevant assets (i.e.models used for the black box), and a logging backend invoked for every oracle observation. To that end, we implement the Protein Objectives Library (poli) to provide potentially isolated black box functions. Building on open source tools, poli currently provides 35 black box tasks; besides the Practical Molecular Optimization (PMO) benchmark (Huang et al.,, 2021; Gao et al.,, 2022; Brown et al.,, 2019), it includes Dockstring (García-Ortegón et al.,, 2022) as well as other protein-related black boxes like stability and solvent accessibility (Delgado et al.,, 2019; Blaabjerg et al.,, 2023; Chapman and Chang,, 2000; Stanton et al.,, 2022).333A complete list can be found in our documentation: https://machinelearninglifescience.github.io/poli-docs/ The majority of black box functions can be queried with any string that complies with the corresponding alphabet making the oracles available for free-form optimization. This is an important distinction compared to pre-existing benchmarks that rely on pools of precompiled observations (Notin et al.,, 2023).
We further provide an interface for the solvers used for the individual optimization tasks: poli-baselines. Consistent, stable (and up to date) environments of individual optimizers can be found therein, as well as a standardized way to query them and solve the problems raised in the previous section. These environments and optimizers are tested weekly through GitHub actions, guaranteeing their usability. An example of how poli and poli-baselines work is provided above. A problem containing a black box, an initial input, and potentially a data package is created through problem factories/a create method. A solver takes as input a black box and optionally supervised data, and uses the solve method to run the optimization for a given number of iterations. Such an interface can accommodate any optimizer that provides a method for running the optimization for a given amount of steps, as well as optimizers that do not accept custom initializations (e.g. Bounce or BAxUS). Sec. A.4 provides an introduction to this software’s technical details, including a description of the logging logic though observers that keep track of every black box call.
[…]
DirectedEvolution | |||||
HillClimbing | |||||
CMAES | |||||
GeneticAlgorithm | |||||
Hvarfner’s VanillaBO | |||||
RandomLineBO | |||||
SAASBO | |||||
Turbo | |||||
BAxUS | |||||
Bounce | |||||
ProbRep |
4.2 Benchmarking HDBO in discrete sequences
Using the black boxes provided in poli, as well as the solvers provided in poli-baselines, we benchmark the performance of high-dimensional Bayesian optimization solvers on discrete sequences. Such optimization can take place either at the sequence level (as solvers in the Structured Spaces do), or in a continuous version via either one-hot representations or learned latent spaces. Benchmarks on both fronts are presented in this paper: from sequence design tasks of varying complexity (mimicking protein engineering) in one-hot/sequence space using Ehrlich functions (Stanton et al.,, 2024), to latent space optimization of small molecules on the Practical Molecular Optimization (PMO) benchmark (Gao et al.,, 2022; Huang et al.,, 2021; Brown et al.,, 2019).444We emphasize that the results provided in this paper are continuously updated on our project’s website.
4.2.1 Sequence design problems of varying complexity
Ehrlich functions (Stanton et al.,, 2024) are closed-form procedurally-generated oracles in which a certain collection of motifs needs to be satisfied in a sequence of a pre-selected length. The oracle’s score is between 0 and 1, and determined by how much of the individual motifs are satisfied. The number of motifs and their length are hyperparameters specified by the user. Sec. A.5.1 provides a detailed introduction. In particular, we test on 4 different configurations: a PestControl equivalent555PestControl is a black box whose input are 25 categorical variables, each of which has 5 potential categories. It is a commonly tested black box in the Structured Spaces literature. with an alphabet size of 5 and a sequence length of 25 (i.e. one motif of length 25). Further, 3 different configurations imitating protein design using the alphabet of 20 amino acids and sequence lengths of 5, 15, and 64 are tested. The motif lengths and number of motifs are (4, 1), (7, 2) and (10, 4) respectively. We initialize PestControlEquiv, and the two small Ehrlich problems with 10 supervised samples, and the latter with 1000, and all methods have an evaluation budget of 300.
We optimize these oracles with representatives from the taxonomy that are frequently tested in the HDBO literature. We select Hvarfner’s VanillaBO, RandomLineBO, Turbo, BAxUS, SAASBO, Bounce, and ProbRep, including also baselines like DirectedEvolution (i.e. greedily mutating the incumbent in sequence space), HillClimbing (which explores the input space by taking random Gaussian steps), and evolutionary algorithms/strategies like GeneticAlgorithm and CMA-ES. Notice that, due to the nature of their implementations, BAxUS and Bounce are not initialized with the same supervised data (but rather according to their implementation: SOBOL sampling in a linear subspace). In this scenario, the continuous solvers are optimizing over one-hot inputs, and the discrete ones work over sequence space. The next section explores latent space optimization.
Table 1 shows the mean best value achieved in the aforementioned 4 problems with one standard deviation for 5 seeds. Problems of low complexity (i.e. PestControlEquiv and Ehrlich with seq. length 5) are readily solvable using one-hot BO, with Bounce solving PestControlEquiv in all 5 replications. Such problems are also easily solvable by naïve baselines like DirectedEvolution. For more complex problems, BO optimizers perform equally or worse than greedy baselines. We hypothesize that naïve baselines will perform worse on black boxes with higher-order effects. Note that SAASBO and ProbRep ran out of memory when instantiating the largest problem.
4.2.2 Benchmarking HDBO on PMO
Hill | Genetic | CMAES | Hvarfner’s | Random | SAASBO | BAxUS | Turbo | Bounce | ProbRep | |
Oracle | Climbing | Algorithm | VanillaBO | LineBO | ||||||
albuterol_similarity | ||||||||||
amlodipine_mpo | ||||||||||
celecoxib_rediscovery | ||||||||||
deco_hop | ||||||||||
drd2_docking | ||||||||||
fexofenadine_mpo | ||||||||||
gsk3_beta | ||||||||||
isomer_c7h8n2o2 | ||||||||||
isomer_c9h10n2o2pf2cl | ||||||||||
jnk3 | ||||||||||
median_1 | ||||||||||
median_2 | ||||||||||
mestranol_similarity | ||||||||||
osimetrinib_mpo | ||||||||||
perindopril_mpo | ||||||||||
ranolazine_mpo | ||||||||||
rdkit_logp | ||||||||||
rdkit_qed | ||||||||||
sa_tdc | ||||||||||
scaffold_hop | ||||||||||
sitagliptin_mpo | ||||||||||
thiothixene_rediscovery | ||||||||||
troglitazone_rediscovery | ||||||||||
valsartan_smarts | ||||||||||
zaleplon_mpo | ||||||||||
Sum (normalized per row) |
Hill | Genetic | CMAES | Hvarfner’s | Random | SAASBO | Turbo | Bounce | ProbRep | |
Oracle | Climbing | Algorithm | VanillaBO | LineBO | |||||
albuterol_similarity | |||||||||
amlodipine_mpo | |||||||||
celecoxib_rediscovery | |||||||||
deco_hop | |||||||||
drd2_docking | |||||||||
fexofenadine_mpo | |||||||||
gsk3_beta | |||||||||
isomer_c7h8n2o2 | |||||||||
isomer_c9h10n2o2pf2cl | |||||||||
jnk3 | |||||||||
median_1 | |||||||||
median_2 | |||||||||
mestranol_similarity | |||||||||
osimetrinib_mpo | |||||||||
perindopril_mpo | |||||||||
ranolazine_mpo | |||||||||
rdkit_logp | |||||||||
rdkit_qed | |||||||||
sa_tdc | |||||||||
scaffold_hop | |||||||||
sitagliptin_mpo | |||||||||
thiothixene_rediscovery | |||||||||
troglitazone_rediscovery | |||||||||
valsartan_smarts | |||||||||
zaleplon_mpo | |||||||||
Sum (normalized per row) |
To benchmark the performance of HDBO on discrete sequences in continuous learned representations, we consider the PMO benchmark (Gao et al.,, 2022; Huang et al.,, 2021; Brown et al.,, 2019).666An introduction to PMO can be found in section A.5.2 We select Hvarfner’s VanillaBO, RandomLineBO, Turbo, BAxUS, SAASBO, Bounce, and ProbRep, including also HillClimbing, CMA-ES and GeneticAlgorithm as baselines. All solvers start from the same initial data to ensure a fair comparison except for BAxUS and Bounce due to the nature of their implementations. We test the aforementioned methods on PMO (Gao et al.,, 2022; Huang et al.,, 2021), which requires a discrete representation of small molecules. Thus, we train two MLP VAEs on SELFIES representations of small molecules using Zinc250k (Irwin et al.,, 2020; Zhu et al.,, 2022). These generative models had 2 and 128 latent dimensions, allowing us to get an impression of how these models scale with dimensionality. We restrict sequences to be of length 70 (adding [nop] tokens for padding); the post-processing renders an alphabet of 64 SELFIES tokens. Details can be found in Sec. A.3.
The average best result over 3 runs (of maximum 300 iterations each) is presented in Tables 2 and 3 for 128D and 2D latent spaces respectively. We see a clear advantage in the optimizers that work on learned representations, instead of in discrete space. Such a discrepancy is to be expected: methods that optimize in latent space have been presented with information prior to their optimization campaigns, while methods like Bounce and ProbRep explore the whole discrete space. Further, the simple baseline is reliably beaten by the continuous alternatives in lower dimensions except Turbo, but this advantage is not as clear in the 128D case, signaling a more complex problem. Some of these tasks, however, are equally challenging for all solvers. deco_hop remains close to the original default value of 0.5, and there is no improvement over valsartan_smarts (which only REINVENT improves on in the original PMO results (Gao et al.,, 2022; Loeffler et al.,, 2024)).
5 Conclusion
In this paper, we surveyed the field of high-dimensional Bayesian optimization (HDBO) focusing on discrete problems. This highlighted the need for (i) a novel taxonomy of the field that emphasizes the differences between methods that rely on unsupervised discrete information, and methods that optimize sequences directly, and (ii) a standardized framework for benchmarking HDBO methods. We approach these in the form of two software tools: poli and poli-baselines. Using these tools, we implemented several HDBO methods and tested them in a standard benchmark for sequence design and small molecule optimization. Our findings suggest that optimizers that work on discrete sequence space (i.e. the structured spaces family) perform as expected on problems of low complexity, but do not scale gracefully to problems with larger dictionaries/higher sequence lengths. In such cases, optimizers that leverage pre-trained latent-variable models have an edge over the other tested methods that work directly on sequence space. That being said, we find that, for low budgets, simpler baselines tend to perform as well or better than most BO methods, echoing recent criticisms of molecular optimization (Tripp and Hernández-Lobato,, 2023). Our framework opens the door to fair and easily-replicable comparisons. We expect poli-baselines to be used by practitioners for running HDBO solvers in up-to-date environments compared across several tasks in our ongoing benchmark, which we plan to expand to other discrete objectives in poli.
Limitations and societal impact.
Although we taxonomize different families of HDBO methods, we have only benchmarked a subset; moreover, the black boxes we use are often simple, closed-form functions or data-driven oracles, limiting the applicability of our benchmark to real-world scenarios. The field of discrete sequence optimization for biology/chemistry needs better black box oracles with more relevance to their respective domains. Another limitation of our approach is the lack of a systematic analysis on the myriad of design choices made on these optimization campaigns; we strived for documenting them properly, but an in-depth analysis is called for in future work. Our benchmark is ongoing and we plan to include further experiments in the project’s website with, hopefully, participation from the community. Finally, we note that optimizing small molecules or proteins opens the door to both drug discovery, but also dual use (Urbina et al.,, 2022).
Acknowledgments and Disclosure of Funding
The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). RM is funded by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). SH was further supported by a research grant (42062) from VILLUM FONDEN as well as funding from the European Research Council (ERC) under the European Union’s Horizon programme (grant agreement 101125993). WB was supported by VILLUM FONDEN (40578). This work was in part supported by the Pioneer Centre for AI (DRNF grant number P1). MGD thanks Sergio Garrido and Anshuk Uppal for feedback on early versions of this document; Peter Mørch Groth for useful discussions; Luigi Nardi, Erik Hellsten and Raffaello Baluyot for feedback on the preprint version of this paper, and Samuel Stanton for feedback on the software.
Appendix A Appendix
A.1 Methods Overview
Method | Date (first occurrence) | Reference | Code Available |
SOLID | January 23, 2019 | Winkel et al., (2021) | ✗ |
Deep GPs | May 7, 2019 | Hebbal et al., (2021) | ✗ |
ASM | July 15, 2017 | Palar and Shimoyama, (2017) | ✗ |
Add-GP-UCB | May 13, 2015 | Han et al., (2021) | ✓ |
TuRBO | December 8, 2019 | Eriksson et al., (2019) | ✓ |
LOL-BO | January 28, 2022 | Maus et al., (2022) | ✓ |
ROBOT | October 20, 2022 | Maus et al., (2023) | ✓ |
REMBO | January 9, 2013 | Wang et al., (2016) | ✓ |
SAASBO | June 10, 2021 | Eriksson and Jankowiak, (2021) | ✓ |
Dropout | August 19, 2017 | Li et al., (2018) | ✗ |
BAxUS | November 28, 2022 | Papenmeier et al., (2022) | ✓ |
LineBO | June 10, 2019 | Kirschner et al., (2019) | ✓ |
ALEBO | December 6, 2020 | Letham et al., (2020) | ✓ |
HeSBO | May 24, 2019 | Nayebi et al., (2019) | ✓ |
BORING | October 5, 2020 | Yenicelik, (2020) | ✗ |
REMBO 2.0 | October 18, 2019 | Binois et al., (2020) | ✗ |
Quantile-GP BO | February 1, 2020 | Moriconi et al., 2020b | ✗ |
Warped REMBO | January 1, 2015 | Binois et al., (2015) | ✗ |
closed-form ASM | September 22, 2020 | Wycoff et al., (2021) | ✓ |
Active manifolds | May 24, 2019 | Bridges et al., (2019) | ✗ |
Deep GPs (MO) | January 1, 2019 | Hebbal et al., (2019) | ✗ |
LADDER | December 6, 2021 | Deshwal and Doppa, (2021) | ✓ |
Attr. Adjustment | August 6, 2018 | Eissman et al., (2018) | ✗ |
VAEs DML | November 1, 2021 | Grosnit et al., (2021) | ✓ |
LSBO | February 28, 2018 | Gómez-Bombarelli et al., (2018) | ✓ |
Weigh. Retraining | October 25, 2020 | Tripp et al., (2020) | ✓ |
MORBO (MO) | September 22, 2021 | Daulton et al., 2022a | ✓ |
TREGO | October 10, 2022 | Diouane et al., (2023) | ✗ |
TRIKE | August 8, 2015 | Regis, (2016) | ✗ |
TRPBO | November 21, 2020 | Zhou et al., (2021) | ✗ |
D-SKIP | December 3, 2018 | Eriksson et al., (2018) | ✓ |
RDUCB | May 29, 2023 | Ziomek and Bou Ammar, (2023) | ✓ |
G-Add-GP-UCB | April 9, 2018 | Rolland et al., (2018) | ✗ |
QFF | December 3, 2018 | Mutny and Krause, (2018) | ✗ |
SI-BO | December 5, 2013 | Djolonga et al., (2013) | ✗ |
SRE-IMGPO | July 9, 2016 | Qian et al., (2016) | ✗ |
HDS | June 27, 2012 | Chen et al., (2012) | ✗ |
AL of LEs | October 24, 2013 | Garnett et al., (2013) | ✓ |
SG-VAE | July 3, 2018 | Lu et al., (2018) | ✗ |
BO+MCMC | April 10, 2017 | Gardner et al., (2017) | ✓ |
Ensemble BO | March 31, 2018 | Wang et al., (2018) | ✓ |
BOCK | March 3, 2018 | Oh et al., (2018) | ✓ |
COMBO | December 8, 2019 | Oh et al., (2019) | ✓ |
MGPC-BO | September 1, 2020 | Moriconi et al., 2020a | ✓ |
G-STAR | December 1, 2016 | Pedrielli and Ng, (2016) | ✗ |
CASMOPOLITAN | June 18, 2021 | Wan et al., (2021) | ✓ |
Vanilla BO | February 25, 2024 | Hvarfner et al., (2024) | ✓ |
Bounce | July 2, 2023 | Papenmeier et al., (2024) | ✓ |
EGP | August 7, 2017 | Rana et al., (2017) | ✗ |
CoBO | December 10, 2023 | Lee et al., (2023) | ✓ |
DSO | February 27, 2024 | Kong et al., (2024) | ✗ |
MPD | January 16, 2023 | Nguyen et al., (2022) | ✓ |
GIBO | November 22, 2021 | Müller et al., (2021) | ✓ |
PR | October 18, 2022 | Daulton et al., 2022b | ✓ |
RPP | May 9, 2016 | Li et al., (2016) | ✗ |
MCTS-VS | October 31, 2022 | Song et al., (2022) | ✓ |
Standard BO | February 5, 2024 | Xu and Zhe, (2024) | ✓ |
Scalable FOBO | June 16, 2022 | Ament and Gomes, (2022) | ✓ |
FOBO | December 6, 2017 | Ahmed et al., (2016) | ✗ |
BOSS | October 2, 2020 | Moss et al., (2020) | ✓ |
SILBO | May 29, 2020 | Chen et al., (2020) | ✓ |
SCoreBO | April 21, 2023 | Hvarfner et al., (2023) | ✓ |
HEBO | December 7, 2020 | Cowen-Rivers et al., (2022) | ✓ |
Tree Add-GP-UCB | May 1, 2021 | Han et al., (2021) | ✓ |
SIR/SDR | July 21, 2019 | Zhang et al., (2019) | ✓ |
GaBO | November 22, 2021 | Jaquier et al., (2020) | ✓ |
ECI | April 18, 2024 | Zhan, (2024) | ✓ |
MAVE-BO | March 8, 2024 | Hu et al., (2024) | ✗ |
CMA-BO | February 5, 2024 | Ngo et al., (2024) | ✓ |
RTDK-BO | October 5, 2023 | Shmakov et al., (2023) | ✗ |
PG-LBO | December 28, 2023 | Chen et al., (2024) | ✓ |
TSBO | May 4, 2023 | Yin et al., (2024) | ✗ |
BODi | March 3, 2023 | Deshwal et al., (2023) | ✓ |
Mahalanobis BatchBO | November 2, 2022 | Horiguchi et al., (2022) | ✓ |
KPCA-BO | April 28, 2022 | Antonov et al., (2022) | ✓ |
PCA-BO | July 2, 2020 | Raponi et al., (2020) | ✓ |
DSA | November 18, 2015 | Ulmasov et al., (2016) | ✓ |
RPA-GP | June 12, 2020 | Delbridge et al., (2020) | ✓ |
HD-GaBO | December 6, 2020 | Jaquier and Rozo, (2020) | ✓ |
Amortized BO | May 27, 2020 | Swersky et al., (2020) | ✗ |
d-KG | December 4, 2017 | Wu et al., (2017) | ✓ |
Prabuchandran’s FOBO | March 1, 2021 | Penubothula et al., (2021) | ✗ |
AlgFOO | March 18, 2021 | Shekhar and Javidi, (2021) | ✗ |
GTBO | October 5, 2023 | Hellsten et al., (2023) | ✓ |
CoCaBo | June 12, 2020 | Ru et al., (2020) | ✓ |
MercBO | February 2, 2019 | Deshwal et al., 2021b | ✓ |
HyBO | July 18, 2021 | Deshwal et al., 2021a | ✓ |
BOCS | July 10, 2018 | Baptista and Poloczek, (2018) | ✓ |
LaMBO | July 22, 2022 | Stanton et al., (2022) | ✓ |
LaMBO-2 | December 12, 2023 | Gruver et al., (2023) | ✓ |
G.M. & Lobato | March 1, 2020 | Garrido-Merchán and Hernández-Lobato, (2020) | ✗ |
BOHAMIANN | December 5, 2016 | Springenberg et al., (2016) | ✓ |
AIBO | February 16, 2023 | Zhao et al., (2024) | ✓ |
CoRel | April 26, 2024 | Michael et al., (2024) | ✓ |
VAE-BO+Inv | July 22, 2022 | Verma et al., (2022) | ✗ |
Uncert | November 09, 2021 | Notin et al., (2021) | ✓ |
GAUCHE | September 21, 2023 | Griffiths et al., (2023) | ✓ |
LLMs | July 21, 2024 | Kristiadi et al., (2024) | ✓ |
AntBO | January 23, 2023 | Khan et al., (2023) | ✓ |
RMF | May 20, 2022 | Kim et al., (2022) | ✗ |
A.1.1 On how HDBO can be applied to discrete sequences
Fig. 4 shows an overview of the different ways in which HDBO could be applied to discrete sequences. On the left, we have the original sequence space with inputs being a list of categorical variables belonging to some alphabet; methods like Bounce, Prob. Rep and those that work on mixed/categorial inputs in the Structured Spaces family (see Sec. 3.7) work in this space. Since the original sequence space is large for the problems we are interested in, several other methods rely on subspaces from the original problem. Such subspaces can be constructed either as linear projections (Sec. 3.4) or as learned non-linear latent variables (Sec. 3.5). Once a latent continuous representation is available, all other families in the taxonomy become available and could be leveraged for latent space optimization. In these cases, there is usually a mapping between the subspace and the original high-dimensional space which allows for computing objective functions. In the linear case an up-projection is given by multiplying with a matrix of the right shape, and in the case of non-linear representations it is usually a decoding process.
In this paper, we explore applying the discrete optimizers of the Structured Spaces family in the sequence spaces of several Ehrlich problems (Sec. 4.2.1), as well as small molecule optimization using SELFIES representations. The continuous optimizers chosen from the other families are applied directly in one-hot space in the case of Ehrlich, or in the latent space of a Variational Autoencoder learned for SELFIES strings (Sec. 4.2.2).
A.2 Reproducing results
Bounce.
We forked the official open source implementation of Bounce777https://github.com/LeoIV/bounce and added an interface between poli’s black boxes and their optimizer. Moreover, we made their implementation pip-installable. Bounce’s implementation is originally provided with an MIT license.
Probabilistic Reparametrization
provides an open-source implementation built on GPyTorch and BoTorch.888https://github.com/facebookresearch/bo_pr They provide a pip installable Python package which did not install until the dependencies mentioned above were fixed (to 1.11 and 0.7 respectively); further, the environment provided had to be updated by replacing the deprecated scikit-learn installation in a fork of their repository. After implementing an interface for poli black boxes, we relied on their script run_one_replication.py to implement a custom solver. The original PR code is provided with an MIT License.
SAASBO, Hvarfner’s Vanilla BO
were all implemented by following the tutorials in Ax. Ax provides models for SAASBO, and we implemented a BoTorch model following the original implementation. We also provide an implementation of ALEBO using Ax. Ax is provided with an MIT License, and Hvarfner’s original code does not have a license in GitHub yet.
Turbo
was implemented by following the tutorial on BoTorch,999https://github.com/pytorch/botorch/blob/main/tutorials/turbo_1.ipynb which can be found on GitHub under MIT License.
BAxUS
is implemented using the Python package provided by the authors.101010https://github.com/LeoIV/BAxUS Since their work builds on the original TuRBO code from Uber, it inherits their license.
A.3 Training VAEs on SELFIES
Models.
We use PyTorch to implement a VAE with [one_hot_input, 1024, 512, 256, latent_dim] encoder and a symmetric decoder, using ReLU activations. The latent space prior used is a standard Gaussian, and the decoded distribution is a categorical. We use torch.distributions to compute ELBO losses without any weighting.111111The exact implementation can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/1a81f7fbd531eb4dc70f84828d6efc1f2ee53e5e/src/hdbo_benchmark/generative_models/vae_selfies.py. We are currently exploring a generative model with more expressive power, and plan to update the benchmarks accordingly in our project’s website.
Training regimes.
We train for a maximum of 1000 epochs using early stopping with a patience of 50 epochs. Our batch sizes are , learning rates are , and optimizer is AdamW.121212The training script can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/master/src/hdbo_benchmark/experiments/training_vae_on_zinc250k/train_vae.py
A.4 Technical details on poli and poli-baselines
This section gives a short overview on the possibilities of poli and poli-baselines. For further information please refer to the full online documentation under https://machinelearninglifescience.github.io/poli-docs/.
The aim of poli is to make it easier to benchmark new algorithms on new problems. There are three components to our framework: problems created by problem factories which contain black boxes, solvers that optimize these problems (aiming always at maximization), and observers which get attached to black boxes and log every single black box call. Fig. 5 shows these three components and their relationships. poli provides a unified numpy (Harris et al.,, 2020) interface, with inputs being numpy arrays of strings, and outputs being numpy arrays of floats. Finally, poli makes sure to log every call to the black box, as well as handling evaluation budgets—function calls are measured transparently and consistently across algorithms.
A major issue on the endeavour of benchmarking on new problems are often conflicting package dependencies. When evaluating a brand-new solver on an older benchmark problem, it can simply be impossible to create a Python environment in which both can run. One may re-implement one or the other, though this is not only cumbersome but also error-prone. With the isolation mechanisms provided by poli this is no longer a problem. In particular, the isolation of the different components in the system (the black box, or the observer that is used for logging) allows an easy comparison to additional algorithms or on other problems, without the need to change plotting scripts.
LABEL:lst:example_experiment shows the general workflow:131313All listings in this appendix were tested using poli version v1.0.1, and poli-baselines version v1.0.2. They can be easily implemented and tested in a Colab notebook. a user instantiates a black box by calling create with the corresponding black box’s name, and the name of an observer. poli then takes care that both black box and observer are instantiated, in different conda environments if necessary. create returns an object inheriting from Problem which provides access to all things necessary: the black box function, initial observations, the observer and problem information like the alphabet and if sequences are aligned or not.141414A future release adds a data package attribute, attached to problems. The data package standardizes the information available to solvers by declaring unsupervised and supervised data; equalizing solver initialization. The black box is a function that takes and returns a numpy array, and the observer is informed automatically of any such calls. Each component of this workflow (black box, algorithm, and observer) can be easily exchanged by user-developed classes more specifically tailored to the need at hand. We refer to our online documentation151515https://machinelearninglifescience.github.io/poli-docs/ for the most recent information on how to do this.
To analyze an experiment a user has to define an observer, inheriting from AbstractObserver. LABEL:lst:example_observer shows how to quickly set up a simple text-file logger. The philosophy is to define the quantities of interest independent of problem or algorithm. We provide standard observers, for example an mlflow observer, and we encourage users to implement their own, to log other metrics of interest. Just as black boxes, observers can run in isolation, such that the environment is independent of problem and, or solver, to facilitate consistent recording of metrics. Whenever observers can run directly in the same environment as the solver, they could be attached directly to the black box using the set_observer method. LABEL:lst:example_observer_without_isolation shows an example using the same workflow as above, and the SimpleObserver presented in LABEL:lst:example_observer.
To expand on the suite of available black boxes, the user has to implement a subclass of the core object of poli: AbstractBlackBox. The method to implement is _black_box, which takes as input an array of strings x (as well as an optional context), and outputs the result of evaluating it as an array of floats. Listing LABEL:lst:example_black_box shows an example code snippet. At run time, another user wanting to test an algorithm just imports the relevant black box object from poli’s repository, and gets an interface to a dynamically instantiated black box potentially running in a different conda environment. The __call__ method then takes care of communication with the caller and logging to an observer. Some of these black boxes require additional assets (e.g. the weights of a neural network, or csv files). They are all either already provided or dynamically downloaded when the black box is used.
Optimization algorithms can be oblivious to any requirements a black box might have. Solvers have only access to the AbstractBlackBox interface. They communicate with the problem only indirectly via a local network socket provided by python’s native multiprocessing library. This is the isolation mechanism that allows both algorithm and problem to run in different python environments.
A.5 Benchmark Introduction
A.5.1 Ehrlich functions
Ehrlich functions (Stanton et al.,, 2024) were proposed as a closed-form alternative to black boxes like FoldX or RaSP, which could potentially raise licensing issues and, before poli, were not readily available for querying. This section explains how Ehrlich functions are constructed and queried from a bird’s eye view. The details can be found in the original paper.
These functions are procedurally generated from a random seed and hyperparameters like alphabet size , sequence length , number of motifs and motif length . The following are the steps that are followed to construct the problem:
1.
A sparse transition matrix of size is constructed. This transition matrix spans a finite Markov chain: sequences in the problem are constructed by starting with a random symbol and following the probabilities determined by . The rows of this transition matrix are probability vectors which determine how likely it is to go from one symbol in the alphabet to another. Such a transition matrix needs to satisfy that (i) the probability of going from one symbol to itself is non-zero, and (ii) it is always possible to go from any symbol to any other symbol after following a sequence of finite steps (irreducibility). In other words, we need the Markov chain spanned by to be ergodic. defines the search space of feasible sequence: the fact that is sparse means that some of the transitions are impossible.
2.
Once has been constructed, motifs of length are sampled according to it by sampling a sequence of length and splitting it. This ensures that the motifs are feasible and easily within reach from one to the next.
3.
These motifs are to be satisfied in certain positions in the string. These positions are constructed using random offsets of size . For example, the motif “ADER” with offsets of is satisfied if the sequence contains “ADXEXR” where “X” can be any other member of the alphabet: “D” is at distance 1 of “A”, “E” is at a distance 3 of “A” and “R” is at distance 5 of “A”. These random offsets are randomly constructed to maximize slackness in the sequence.
We take the original implementation161616https://github.com/prescient-design/holo-bench and wrap our black box logic around it, making it compatible with solvers in poli-baselines. We also include a data package made by using the original utilities for sampling the underlying transition matrix.
A.5.2 PMO
The Practical Molecular Optimization (PMO) benchmark contains a representative set of tasks defined on small molecule inputs with respective computational oracles. The input to the black box functions are alphabet representation for small molecules as either tokenized SMILES (Weininger,, 1988) or SELFIES (Krenn et al.,, 2020) (see below) for the exact alphabets used – in principle one can be converted into the other. While small molecule sequences usually contain fewer elements in a sequence than for example proteins, the alphabet can contain more tokens making this yet another high dimensional discrete optimization space. We build upon the work by Gao et al., (2022), who propose molecular optimization focused on validity, diversity, synthesizability - using computational values for all metrics and a budget of 10.000 evaluations. The PMO suite itself is based on the existing benchmarks Guacamol (Brown et al.,, 2019), and elements of the TDC (Huang et al.,, 2021). The types of optimization tasks for small molecules can be differentiated into: optimizing for simple metrics like QED, LogP171717These are aggregate properties of small molecules, and can sometimes be poor proxies for other chemical downstream behavior. qualitative tasks, such as scaffold hopping, rediscovery of particular substances (e.g. troglitazone), and more complex tasks such as docking surrogates and classifiers on fingerprint representations of molecules (GSK3, JNK3, DRD2) (Huang et al.,, 2021). We evaluate a selection of solvers across all tasks with multiple seeded runs on a fixed budget and report the average of the best observations, weighing all functions equally.181818Equal weighting can favor methods that perform well on a few exploitable tasks (e.g. logP, QED). We draw the users attention to that fact with the alternatives to discount or discard such tasks. Altogether, these tasks constitute a set of functions which can be optimized and which have previously been used to assess algorithm performance in the bio-chemical domain, allowing a comparisons to previous results and reported benchmarks. We make these environments accessible and solvable in a unified way through the poli infrastructure.
A.6 Compute details for all experiments
A.6.1 Ehrlich functions
The following solvers ran in an HPC with memory/CPU/GPU requirements given by 30G/10/1 Titan X with 12Gb of memory with: SAASBO, ProbRep, Hvarfner’s VanillaBO, RandomLineBO, DirectedEvolution, HillClimbing, CMA-ES and GeneticAlgorithm. All these were time-gated via SLURM to run for at most 24hrs.
The remaining solvers (BAxUS, Turbo and Bounce) ran in a Google Cloud instance with one Tesla T4 with 16Gb.
A.6.2 PMO
HillClimbing, Hvarfner’s VanillaBO, RandomLineBO, SAASBO and ProbRep experiments ran in an HPC cluster on CPUs using equivalent SLURM scripts (max 24h of runtime).
Turbo ran on an M2 Max Mac with 32Gb of memory using MPS.
BAxUS and Bounce ran on a Deep Learning compute server using the marketplace solutions of Google Cloud Platform, using a Tesla T4 (approx. 16Gb of memory).
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] MGD: yes, RM: yes, SB: yes, YZ: yes, SH: yes, WB: yes
Did you describe the limitations of your work? [Yes] MGD: yes, RM: yes, SB: yes, YZ: yes, SH: yes, WB: yes
Did you discuss any potential negative societal impacts of your work? [Yes] MGD: yes, RM: yes, SB: yes, YZ: yes, SH: yes, WB: yes
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] MGD: yes, RM: yes, SB: yes, YZ: yes, SH: yes, WB: yes
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A] No theoretical results.
Did you include complete proofs of all theoretical results? [N/A] No theoretical results.
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the project URL you can find a link to our repository with instructions.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Sec. A.3.
- (c)
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Sec. A.6 in the appendix
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] We use the PMO benchmark, for which we cite not only the original developers, but also the work they are based on plus their TDC framework. Plus, we cite the authors of every HDBO solver we could find, let alone use. For Ehrlich functions, we cite the relevant paper.
Did you mention the license of the assets? [Yes] Yes, see Sec. A.2 in the appendix.
Did you include any new assets either in the supplemental material or as a URL? [Yes] In the project website. https://machinelearninglifescience.github.io/hdbo_benchmark/
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] We are not using data on people.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] We are not using data on people.
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We are not crowdsourcing experiments, nor conducting research on humans.
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] We are not crowdsourcing experiments, nor conducting research on humans.
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] We are not crowdsourcing experiments, nor conducting research on humans.